OpenStudy (anonymous):

The sum of two numbers is 86 and the difference is 34, what are the numbers?

OpenStudy (zzr0ck3r):

assume b>a a+b=86 b-a=34

OpenStudy (zzr0ck3r):

we may assume a is not b, because we have b-a=34

OpenStudy (zzr0ck3r):

then we may assume without loss of generality that one is bigger than the other, i pick b

OpenStudy (anonymous):

I am looking for two numbers.

OpenStudy (zzr0ck3r):

right, we now have a system of equations you need to solve a+b=86 b-a=34

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

does this seem familiar at all? what are you doing in class?

OpenStudy (anonymous):

26 and 60?

OpenStudy (zzr0ck3r):

yep

OpenStudy (anonymous):

okay help with a few more?

OpenStudy (anonymous):

wait nvm

OpenStudy (mathmate):

@blveXO For future reference, sum and difference problems can be solved as follows: big number = (sum+difference)/2 = (86+34)/2=60 small number = (sum - difference)/2 = (86-34) /2 = 26